One recipe for granola makes six servings and requires 2/3 cup of sugar. A similar recipe makes eight servings and calls for 3/4 cup of sugar. Use comparison of rational numbers to determine which recipe contains more sugar per serving.

Please Help I do not know who to start.

Answer by ewatrrr(24785)


Hi,
(2/3)c/6 servings = (1/9)c per serving Or .11c per serving
(3/4)c/8 servings = (3/32)c per serving Or .09c per serving
(2/3)c/6 servings contains more sugar per serving.